According to the U.S. Energy Information Administration (EIA), Americans consumed 140.43 billion gallons ofgasoline in 2015, or about 384.74 million gallons each day. It’s generally thought that one can produce on average 7.5 metric tons of corn per hectare, and achieve about a 30% yield of ethanol from this corn (i.e., 0.3 kg ethanol/kg corn). The total amount of land in all active farms in the U.S. is 415 million acres. (a) How much farmland would be required to replace the current volume of gasoline with ethanol? (b) How does that compare to the total amount of land now being farmed? (Hint: In working this problem, be careful with units, particularly hectares and acres. Comment: Note that your answer to (a) relates simply to volumes of the two liquids, not replacing the amount of energy in gasoline with energy from ethanol.)
Answer #1

Part a

Volume of gasoline required Vg = 384.74*10^6 gal/day

Volume of ethanol required = Vg = 384.74*10^6 gal/day

Density of ethanol = 6.585 lb/gal

Mass of ethanol required = volume x density

= (384.74*10^6 gal/day) x (6.585 lb/gal)

= 2.533 x 10^9 lb/day x 1kg/2.205lb

= 1.149 x 10^9 kg/day

1 kg corn produces = 0.3 kg ethanol

Mass of corn required

= (1 kg corn/0.3 kg ethanol) x (1.149 x 10^9 kg ethanol/day)

= (3.83 x 10^9 kg corn/day) x (1MT/1000kg)

= 3.83 x 10^6 MT corn/day

Land required

= (1 hectare / 7.5 MT corn) x (3.83 x 10^6 MT corn/day)

= 5.11 x 10^5 hectare

Part b

total amount of land in all active farms in the U.S.

= 415 * 10^6 acre x 1hectare/2.471acre

= 1.679 x 10^8 hectare

Percent of land in corn farming

= 5.11 x 10^5*100/ 1.679 x 10^8

= 0.30%